# von Neumann entropy and quantum version of thermodynamic entropy

### v2

@@ III.2 As entropy of thermodynamic ensemble
As in Sec. II.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero (see Eq. 1). In this case, as we see from Eq. 3, the VN entropy is also zero. When the system is not in pure state, rather in a statistical mixture of pure states (represented by $\rho$), a formula for entropy must quantify the amount of “mixedness” (or statistical impurity) in the state $\rho$, similar to how the BN entropy quantifies the amount of “uncertainty” of a system to be found in a microstate. If we now see the decomposition in Eq. 4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-k_{B}\sum_{i}p_{i}\ln p_{i}$, which is the Shannon (SN) entropy for the probability distribution $\{p_{i}\}$. Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable. One way to circumvent this problem is to consider all possible decompositions of $\rho$ and take the one for which $S_{SN}(\{p_{i}\})$ is minimum. The advantage of defining entropy this way is that the entropy is now unique for a given $\rho$. In fact it can be proved that the VN entropy of a given $\rho$ corresponds to the aforementioned minimum [17, 18]:
@@
Another way to understand the VN entropy is as follows. If we take a set of orthonormal states to be a set of distinguishable microstates, we can define an entropy from the probabilities associated with these microstates. If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-k_{B}\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. Unfortunately, again, this entropy is not unique as it depends on the set of orthonormal states (microstates). To understand how to avoid this problem, we first define a new density matrix $\rho^{\prime}=\sum_{n}w_{n}\ket{\phi_{n}}\bra{\phi_{n}}$ |, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. We may note that |$\ket{\phi_{n}}$ is an eigenket of $\rho^{\prime}$ with eigenvalue $w_{n}$. This density matrix ($\rho^{\prime}$) is in the same form as the one defined in Eq. 5. For this updated density matrix after measurement ($\rho\to\rho^{\prime}$), as mentioned earlier, von Neumann showed that $-\text{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\text{Tr}(\rho\ln\rho)$ [8]. Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$. Accordingly,
@@
III.3 Addressing objections against VN entropy
@@
The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves (unitarily) with time. As seen earlier, $S_{VN}(t)=S_{VN}(0)$ when $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$. It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium. Does not the VN entropy apply to systems in non-equilibrium?
At the outset, we stress that the above issue arises only when the system is in a non-equilibrium state and we want to know the entropy during a process. In the thermal equilibrium, as discussed earlier in Sec. III.1, the VN entropy and the TH entropy are equivalent (under certain general conditions). So, the VN entropy is capable of determining the change in (TH) entropy due to a process (the system is in equilibrium before and after the process).
@@
While trying to resolve the time-invariance problem of the VN entropy, one can think of two major different approaches. First, one considers that the VN entropy formula needs modification (or some kind of generalization) so that the entropy from the new formula increases (and saturates in equilibrium) as the system evolves with time. The observational entropy is one such popular alternative to the VN entropy [16, 21, 22]. Second, instead of considering that the VN entropy formula is wrong, one may consider that the time evaluation of the density matrix $\rho$ is not correctly described by the von Neumann-Liouville equation. We prefer the second option for the following reasons.
@@
The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored). The subadditivity property is stated as follows: if a bipartite system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices of the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [26]. This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case). This contradicts the notion of the additivity property that the TH entropy has. We can circumvent this apparent contradiction in the following way.
@@
With the above understanding, we now check if the VN entropy of a subsystem has the additivity property. It is known that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [30, 31, 32, 33]. In fact, for our system with $\Delta_{2}=0.5$ (see Sec. IV.3), we see that the average VN entropy $\overline{S}^{sb}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size; relevant results can be found in Fig. 1.
Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}^{sb}_{VN}$) grows linearly with subsystem size. Here total system size is $N=16$ and $\Delta_{2}=0.5$.
@@
IV Entropy of large quantum systems
In Sec. III.1, we have seen that the VN entropy of a thermalized isolated or open system is a basis-independent representation of the quantum statistical entropy (which, under certain conditions, is equivalent to the TH entropy). There we start with the qBN or qGB entropy and then show that it is simply equal to the VN entropy corresponding to some density matrix (we were not much concerned about any significance or possible interpretation of the density matrix).
@@ IV Entropy of large quantum systems
In this section, we go in the opposite direction: we first calculate the VN entropy independently and then verify if that is equal (equivalent) to the TH entropy of the system (subsystem). To calculate the VN entropy, we take an appropriate density matrix which describes the system (subsystem) in thermodynamic equilibrium.
In the following, we discuss the issue of equivalence separately for the isolated, open and subsystem (of a much larger system).
IV.1 Isolated quantum system
@@
We consider an isolated quantum system in thermal equilibrium. As discussed in Sec. II.3, a large isolated quantum system will be in a mixed state, for one or multiple reasons. An isolated system in thermal equilibrium is appropriately described by a microcanonical (mc) ensemble. Assume that we prepare the system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let {|$\{\ket{n}\}$ be the set of eigenkets of the Hamiltonian of the system, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$. The density matrix which describes the isolated system is given by,
$\rho_{mc}=\sum_{\begin{subarray}{c}n\\ E_{n}\in\left(E,E+\Delta E\right]\end{subarray}}p_{n}\ket{n}\bra{n},$ (8)
@@
where $p_{n}$ is the probability to find the system in state |$\ket{n}$. Since $\Delta E$ is macroscopically small, all the accessible microstates (here, the eigenkets) will essentially have almost the same energy. Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{D}$ (see aslo Sec. II.3). As $S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})$, we need the eigenvalues of $\rho_{mc}$ to efficiently compute $S_{VN}$. It is clear from Eq. 8 that |$\ket{n}$ is an eigenket of $\rho_{mc}$ with eigenvalue $\frac{1}{D}$, i.e., $\rho_{mc}\ket{n}=\frac{1}{D}\ket{n}$ for any energy eigenket from the shell. This implies that
$S_{VN}=-k_{B}\text{Tr}(\rho_{mc}\ln\rho_{mc})=k_{B}\ln D.$ (9)
If we now directly calculate the TH entropy (or equivalently, the qBN entropy), following our discussion in Sec. II.3, it would be
$S_{TH}\equiv S_{qBN}=k_{B}\ln D,$ (10)
where $D$ is the number of eigenvalues within the shell. It may be noted here that one can subdivide the energy shell into different macrostates (subspaces), with one of them having the largest number of microststes (the largest dimension) [7, 5]. For a thermodynamically large system, the dimension of this largest subspace, which corresponds to the thermal equilibrium, is essentially equal to the dimension of the shell ($D$).
@@
We see from Eqs. 9 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium. The detailed discussion on the conditions under which this equivalence works can be found in Sec. II.2.
IV.2 Weakly interacting open quantum system
@@
As discussed in Sec. II.3, the density matrix that appropriately describes the open system is $\rho_{\beta}=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy. In fact, if the system-bath combination is described by a microcanonical ensemble or for the overwhelmingly large number of pure states (of the system-bath combination), it can be quite generally shown that an open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [11, 10]. Accordingly, the VN entropy of the open system is,
$S_{VN}=-k_{B}\text{Tr}(\rho_{\beta}\ln\rho_{\beta})=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (11)
@@
Figure 2: The subsystem VN entropy ($S^{sb}_{VN}$) for the individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system. Here total system size is 16 and subsystem size is 6.
As discussed in Sec. II, for the open quantum system, the TH entropy (or equivalently, the qGB entropy) is given by:
$S_{TH}\equiv S_{qGB}=-k_{B}\sum_{n}p_{n}\ln{p_{n}},$ (12)
where the expression of $p_{n}$ is the same as the one appears in Eq. 11.
Figure 3: The average VN entropy ($\overline{S}^{sb}_{VN}$) of subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum). Here, the system (subsystem) size is 16 (6) and $\Delta_{2}=0.5$.
@@
We now see from Eqs. 11 and 12 that the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium. The detailed discussion on the conditions under which this equivalence works can be found in Sec. II.2.
IV.3 Subsystem of a quantum system
@@
We separately discuss here the question of equivalence between the subsystem entropies when the full system is in a mixed state (described by a microcanonical ensemble) or in a pure state (both typical and atypical).
@@
If the full system is in a mixed state $\rho_{mc}$, described by a microcanonical (mc) ensemble, then the subsystem will be in a Gibbsian distribution $\rho_{\beta}$ [11, 10]. If $\rho_{sb}$ is the reduced density matrix for the subsystem, i.e., $\rho_{sb}=\text{Tr}_{B}(\rho_{mc})$, then $\rho_{sb}\approx\rho_{\beta}=\frac{e^{-\beta H}}{Q}$ with $Q=\text{Tr}_{sb}(e^{-\beta H})$. It may be noted that the first trace (Tr) operation here is over the degrees of freedom of the bath, B (complement of the subsystem in the full system), and the second trace operation is over the degrees of freedom of the subsystem (sb). We also note that, here, $H$ is the Hamiltonian of the subsystem if the interaction between the subsystem and the bath can be ignored, else it is the effective Hamiltonian of the subsystem, determined by $H=-\frac{1}{\beta}\ln(Q\text{Tr}_{B}(\rho_{mc}))$. This equation can be solved self-consistently: initially, we calculate $Q$ from the Hamiltonian of the subsystem (without interaction with the bath) and get the first order approximation of the effective $H$. Next we use this effective $H$ in $Q$ to find the next order approximation of the effective $H$. We continue this till we get a self-consistent result. Fortunately, we do not need to calculate this effective $H$ in the numerical calculations that follow.
@@
If the full system is in a pure state, for the overwhelming majority of pure states (of the total system), the subsystem can be described by a Gabssian distribution $\rho_{\beta}$ [11, 10]. This important result, known as canonical typicality, leaves open the possibility of pure states (of the full system) for which a subsystem is not described by a Gibbsian distribution or for which a subsystem does not thermalize. We note that there are systems whose eigenstates (from a narrow energy window) are such that the reduced density matrices calculated from them are quite different to each other. These systems do not follow the eigenstate thermalization hypothesis or ETH [27, 28, 5].
Above important results immediately imply the following: if the full system is described by a microcanonical ensemble, the VN entropy of a subsystem is equivalent to its TH entropy (qGB entropy). If the full system is in a pure state, for the overwhelming cases (i.e. for the typical states), the VN entropy of a subsystem will be equivalent to its TH entropy (qGB entropy). But for systems for which the ETH does not hold, this equivalence will not work in general (corresponding pure states are atypical). In the following we take a spin system and numerically demonstrate the issue of equivalence between the VN entropy and TH entropy of a subsystem in the different cases mentioned above.

### v3

@@ 3.1 Entropy of thermodynamic ensemble
As in Sec. 2.3, if a pure quantum state is taken to be a microstate, then the qBN entropy of a system in a pure quantum state is zero (see Eq. 1). In this case, as we see from Eq. 3, the VN entropy is also zero. When the system is not in pure state, rather in a statistical mixture of pure states (represented by $\rho$), a formula for entropy must quantify the amount of “mixedness” (or statistical impurity) in the state $\rho$, similar to how the BN entropy quantifies the amount of “uncertainty” of a system to be found in a microstate. If we now see the decomposition in Eq. 4, it would suggest that the corresponding entropy of the system is $S_{SN}(\{p_{i}\})=-\sum_{i}p_{i}\ln p_{i}$, which is the Shannon (SN) entropy for the probability distribution $\{p_{i}\}$. Unfortunately, the decomposition of a mixed state is not unique, hence this way of defining entropy is unacceptable. One way to circumvent this problem is to consider all possible decompositions of $\rho$ and take the one for which $S_{SN}(\{p_{i}\})$ is minimum. The advantage of defining entropy this way is that the entropy is now unique for a given $\rho$. In fact it can be proved that the VN entropy of a given $\rho$ corresponds to the aforementioned minimum [15, 8]:
@@
Another way to understand the VN entropy is as follows. If we take a set of orthonormal states to be a set of distinguishable microstates, we can define an entropy from the probabilities associated with these microstates. If {|$\{\ket{\phi_{n}}\}$ is a set of orthonormal states, the entropy one can define is $S_{SN}(\{w_{i}\})=-\sum_{n}w_{n}\ln w_{n}$, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. Unfortunately, again, this entropy is not unique as it depends on the set of orthonormal states (microstates). To understand how to avoid this problem, we first define a new density matrix $\rho^{\prime}=\sum_{n}w_{n}\ket{\phi_{n}}\bra{\phi_{n}}$ |, where $w_{n}=\bra{\phi_{n}}\rho\ket{\phi_{n}}$. We may note that |$\ket{\phi_{n}}$ is an eigenket of $\rho^{\prime}$ with eigenvalue $w_{n}$. This density matrix ($\rho^{\prime}$) is in the same form as the one defined in Eq. 5. For this updated density matrix after measurement ($\rho\to\rho^{\prime}$), as mentioned earlier, von Neumann showed that $-\mathrm{Tr}(\rho^{\prime}\ln\rho^{\prime})\geq-\mathrm{Tr}(\rho\ln\rho)$ [36]. Thus we see that the VN entropy corresponds to the minimum of $S_{SN}(\{w_{i}\})$, which we obtain when we choose the microstates as the eigenkets of the given density matrix $\rho$. Accordingly,
@@
3.2 Addressing objections against VN entropy
@@
The first of these objections is that the VN entropy of an isolated quantum system does not change (more specifically, increase) as the system evolves (unitarily) with time. As seen earlier, $S_{VN}(t)=S_{VN}(0)$ when $\rho(0)\to\rho(t)$ under the time evaluation operator $U(t)$. It is known that, for an isolated system, the TH entropy increases as the system approaches thermal equilibrium. Does not the VN entropy apply to quantum systems in non-equilibrium?
At the outset, we stress that the above issue arises only when the system is in a non-equilibrium state and we want to know the entropy during a process. In the thermal equilibrium, as discussed earlier in Sec. 2 and analyzed in details in Sec. 4, the VN entropy and the TH entropy are equivalent (under certain general conditions). So, the VN entropy is capable of determining the change in (TH) entropy due to a process (the system is in equilibrium before and after the process).
@@
While trying to resolve the time-invariance problem of the VN entropy, one can think of two major different approaches. First, one considers that the VN entropy formula needs modification (or some kind of generalization) so that the entropy from the new formula increases (and saturates in equilibrium) as the system evolves with time. The observational entropy is one such popular alternative to the VN entropy [35, 37, 21]. Second, instead of considering that the VN entropy formula needs modification, one may consider that the time evaluation of the density matrix $\rho$ is not correctly described by the von Neumann-Liouville equation. We prefer the second option for the following reasons.
@@
The second major objection against the VN entropy is that it obeys the subadditivity property while the TH entropy is additive (for a large uniform system where the interaction energy between its two parts can be ignored). The subadditivity property is stated as follows: if a bipartite system ($AB$) is in state $\rho_{AB}$, and its parts $A$ and $B$ are respectively in states $\rho_{A}$ and $\rho_{B}$ (which are reduced density matrices from the state $\rho_{AB}$), then $S_{VN}(\rho_{AB})\leq S_{VN}(\rho_{A})+S_{VN}(\rho_{B})$ [2]. This implies that, if the system is in a pure state, its VN entropy will be zero but the sum of the VN entropies of its parts is positive (the sum can be zero in a very special case). This contradicts the notion of the additivity property that the TH entropy has. We can circumvent this apparent contradiction in the following way.
@@
With the above understanding, we now check if the VN entropy of a subsystem has the additivity property. It is known that the VN entropy follows a volume law for excited states (especially the states from the middle of its spectrum) [18, 1, 20, 3]. In fact, for our system with $\Delta_{2}=0.5$ (Sec. 4.4), we see that the average VN entropy $\overline{S}^{sb}_{VN}$ (averaged over all the states from a narrow energy window in the middle of the spectrum) is extensive and grows linearly with subsystem size; relevant results can be found in Fig. 1.
Figure 1: In the middle of the spectrum, the average VN entropy ($\overline{S}^{sb}_{VN}$) grows linearly with subsystem size. The total system is described by the Hamiltonian of Eq. 20 and $\overline{S}^{sb}_{VN}$ is calculated as per Eq. 4.4. The calculations are done with the total system size $N=16$ and $\Delta_{2}=0.5$.
@@
4 Entropy of large quantum systems
@@
In this section, we separately calculate the VN entropy and the TH entropy (represented by qBN or qGB entropy) of the system (or subsystem), and compare if they are equal or equivalent. To calculate the VN entropy, we take an appropriate density matrix which describes the system (subsystem) in thermodynamic equilibrium. In the following, we discuss the issue of equivalence separately for the isolated, open, and subsystem.
4.1 Isolated quantum system
@@
We consider an isolated quantum system in thermal equilibrium. As discussed in Sec. 2.3, a large isolated quantum system will be in a mixed state, for one or multiple reasons. An isolated system in thermal equilibrium is appropriately described by a microcanonical (mc) ensemble. Assume that we prepare the system in such a way that its energy lies in the narrow energy shell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [19]. Let {|$\{\ket{n}\}$ be the set of eigenkets of the Hamiltonian of the system, and $E_{n}$ be the eigenvalue corresponding to |$\ket{n}$. The density matrix which describes the isolated system is given by,
$\rho_{mc}={\sum_{n}}^{\prime}p_{n}\ket{n}\bra{n},$ (8)
@@
where $p_{n}$ is the probability to find the system in state |$\ket{n}$ and the primed sum is for the states associated with the sell $(E-\frac{\Delta E}{2},E+\frac{\Delta E}{2}]$. Since $\Delta E$ is macroscopically small, all the accessible microstates (here, the eigenkets) will essentially have almost the same energy. Application of the principle of equal a priori probabilities suggests that $p_{n}=\frac{1}{d_{E}}$ (see aslo Sec. 2.3). The VN entropy of the system is now given by $S_{VN}=-\mathrm{Tr}(\rho_{mc}\ln\rho_{mc})$. One needs the eigenvalues of $\rho_{mc}$ to efficiently compute $S_{VN}$. It is clear from Eq. 8 that any energy eigenket |$\ket{n}$ is an eigenket of $\rho_{mc}$ with eigenvalue $\frac{1}{d_{E}}$, i.e., $\rho_{mc}\ket{n}=\frac{1}{d_{E}}\ket{n}$. This implies that
$S_{VN}=-k_{B}\mathrm{Tr}(\rho_{mc}\ln\rho_{mc})=\ln d_{E}.$ (9)
If we now directly calculate the TH entropy (or equivalently, the qBN entropy), following our discussion in Sec. 2.3, it would be
$S_{TH}\equiv S_{qBN}=\ln d_{E},$ (10)
where $d_{E}$ is the number of eigenvalues within the shell.
@@
Although $S_{VN}$ and $S_{TH}$ are here calculated separately, we see from Eqs. 9 and 10 that, for an isolated quantum system, the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$), provided that the system is large and in thermal equilibrium. The detailed discussion on the conditions under which this equivalence works can be found in Secs. 2.2 and 2.3.
4.2 Weakly interacting open quantum system
@@
If the system-bath combination is described by a microcanonical ensemble or for the overwhelmingly large number of pure states (of the system-bath combination), it can be quite generally shown that an open quantum system is very well described by a (reduced) density matrix which is equivalent to the known Gibbsian distribution [26, 11]. Denoting the density matrix by $\rho_{\beta}$, we have $\rho_{\beta}=\frac{e^{-\beta H}}{Tr(e^{-\beta H})}$, where the inverse temperature can be fixed by solving the equation $Tr(He^{-\beta H})=U$ with $U$ being the (given) internal or average energy. Accordingly, the VN entropy of the open system is,
$S_{VN}=-\mathrm{Tr}(\rho_{\beta}\ln\rho_{\beta})=-\sum_{n}p_{n}\ln{p_{n}},$ (11)
@@
As discussed in Sec. 2, for the open quantum system, the TH entropy (or equivalently, the qGB entropy) is given by:
$S_{TH}\equiv S_{qGB}=-\sum_{n}p_{n}\ln{p_{n}},$ (12)
where the expression of $p_{n}$ is the same as for the one appears in Eq. 11.
@@
We now see from Eqs. 11 and 12 that the VN entropy is the same as the TH entropy ($S_{VN}=S_{qBN}\equiv S_{TH}$) for a weakly interacting open quantum system in thermal equilibrium. The detailed discussion on the conditions under which this equivalence works can be found in Secs. 2.2 and 2.3.
4.3 Subsystem of a quantum system
@@
We now discuss the question of equivalence between the VN entropy and TH entropy of a subsystem. We separately discuss two different cases - when the full system is in a mixed state (described by a microcanonical ensemble) and when the full system is in a pure state. In the following, the full system is assumed to have two parts - the smaller part is called the “subsystem” (sb) and the rest of the system is called the “bath” (B).
@@
If the full system is in a mixed state $\rho_{mc}$, described by a microcanonical (mc) ensemble, then the subsystem will be in a Gibbsian distribution $\rho_{\beta}$ [26, 11]. If $\rho_{sb}$ is the reduced density matrix for the subsystem, i.e., $\rho_{sb}=\mathrm{Tr}_{B}(\rho_{mc})$, then $\rho_{sb}\approx\rho_{\beta}=\frac{e^{-\beta H}}{Q}$ with $Q=\mathrm{Tr}_{sb}(e^{-\beta H})$. It may be noted that the first trace (TrB) operation here is over the degrees of freedom of the bath (B), and the second trace operation (Trsb) is over the degrees of freedom of the subsystem (sb). We also note that, here, $H$ is the Hamiltonian of the subsystem and $\beta$ is the effective temperature determined by solving the following equation: $\lambda U=\mathrm{Tr}(H\rho_{\beta})$. Here $U$ is the given (fixed) internal energy of the full system and $\lambda$ is the ratio of the subsystem size to the full system size. Now following our discussion for the open system (see Sec. 4.2), we can conclude that if the full system is described by a microcanonical ensemble, the VN entropy of a subsystem is equivalent to its TH entropy (i.e., qGB entropy). We also show this equivalence numerically in the next subsection (Sec. 4.4). For this numerical verification, we independently calculate the VN entropy (as per Eq. 4.4) and the TH entropy (as $\ln(DOS)$; see Eq. 21) of the subsystem. We find that these two are linearly proportional to each other (as in Fig. 3), confirming their equivalence (also see Sec. 4.4).
We now discuss when and how the equivalence between the VN entropy and the TH entropy of a subsystem holds if the full system is in a pure state. We take two different approaches here. In the first approach, we start with the landmark results of Page [22] and Sen [30]. In the second approach, we appeal to the concept of canonical typicality [26, 11].
If a quantum system of Hilbert space dimension $D_{1}D_{2}$ is in a random pure state, then the average VN entropy of the subsystem of Hilbert space dimension $D_{1}$ is
$\mathbb{E}({S}^{sb}_{VN})=\left(\sum_{k=D_{2}+1}^{D_{1}D_{2}}\frac{1}{k}\right)-\frac{D_{1}-1}{2D_{2}}.$ (13)
In the thermodynamic limit (with $1\ll D_{1}\ll D_{2}$), this result, first conjectured by Page [22] and later proven by Sen [30], becomes
$\mathbb{E}({S}^{sb}_{VN})=\ln(D_{1})+O(D_{1}/D_{2}).$ (14)
We argue here that, for a quantum system obeying ETH, the result of Eq. 14 can be strengthened, and further argue that this stronger result implies equivalence of the VN entropy and the TH entropy for a subsystem.
Consider the subspace spanned by the eigenkets of the total system corresponding to a narrow energy shell: $\mathcal{H}_{E}=\mathrm{Span}\{\ket{n}|E_{n}\in(E+\frac{\Delta E}{2},E-\frac{\Delta E}{2}]\}$, where $E_{n}$ is the eigenvalue corresponding to the eigenket |$\ket{n}$ of the total system. The shell width $\Delta E$, as mentioned earlier, is arbitrary, but is microscopically small and microscopically large. Let |$\ket{n_{1}},\ket{n_{2}}\in\mathcal{H}_{E}$, and corresponding reduced density matrices for the subsystem are $\rho_{1}=\mathrm{Tr}_{B}(\ket{n_{1}}\bra{n_{1}})$ and $\rho_{2}=\mathrm{Tr}_{B}(\ket{n_{2}}\bra{n_{2}})$. According to ETH, $\mathrm{Tr}_{sb}(\rho_{1}A)\approx\mathrm{Tr}_{sb}(\rho_{2}A)$, where $A$ is any local operator associated with the subsystem (sb). This hypothesis implies that the sets of eigenvalues of $\rho_{1}$ and $\rho_{2}$ can not be much different, and consequently, $S^{sb}_{VN}(\rho_{1})\approx S^{sb}_{VN}(\rho_{2})$, where $S^{sb}_{VN}(\rho):=-\mathrm{Tr}_{sb}(\rho\ln\rho)$ is the VN entropy of the subsystem corresponding to the reduced density matrix $\rho$. This result is also evident from Fig. 2.
Figure 2: The subsystem VN entropy ($S^{sb}_{VN}$) for the individual eigenkets and $\ln(DOS)$ are plotted against the energy of the full system (described by the Hamiltonian of Eq. 20). Here total system size is 16 and subsystem size is 6.
When we randomly choose states (of the full system) without restriction on the energy scale $E$, then all the energy scales associated with the subsystem are important. Accordingly, all states ($D_{1}$ in number) of the subsystem would contribute to its average VN entropy. This is an intuitive way to understand the result in Eq. 14. When attention is restricted to a pure states in the subspace $\mathcal{H}_{E}$ associated with a narrow energy shell, all the states of the subsystem are no more equally significant. If $l_{1}$ ($l_{2}$) is the size of the subsystem (bath), then the states of the subsystem corresponding to the energy scale $\frac{l_{1}}{l_{1}+l_{2}}E$ would be important. Let $\widetilde{d}_{1}$ is the effective number of states of the subsystem corresponding to its energy scale $\frac{l_{1}}{l_{1}+l_{2}}E$. We can also view $\widetilde{d}_{1}$ as the dimension of the subsystem’s effective Hilbert space associated with the subspace $\mathcal{H}_{E}$. We then expect:
$\overline{S}^{sb}_{VN}\approx\ln\widetilde{d}_{1},$ (15)
where the average is over all eigenkets from the subspace $\mathcal{H}_{E}$. To estimate $\widetilde{d}_{1}$, we note that
$\widetilde{d}_{1}\widetilde{d}_{2}=d_{E},$ (16)
where $\widetilde{d}_{2}$ is similarly defined dimension of the effective Hilbert space of the bath (B) associated with the subspace $\mathcal{H}_{E}$ and $d_{E}$ is the dimension of the subspace, i.e., $d_{E}:=\dim{(\mathcal{H}_{E})}$. We now define a parameter $\eta$ (may be called the effective Hilbert space dimension of a single spin) so that $\eta^{l_{1}}:=\widetilde{d}_{1}$ and $\eta^{l_{2}}:=\widetilde{d}_{2}$. This is similar to $d^{l_{1}}=D_{1}$ and $d^{l_{2}}=D_{2}$ where $d$ is the Hilber space dimension of a single spin (for spin-1/2 system, $d=2$). This shows that
$\frac{\ln\widetilde{d}_{1}}{\ln\widetilde{d}_{2}}=\frac{\ln D_{1}}{\ln D_{2}}.$ (17)
Now starting from Eq. 16, and using Eq. 17, we get, $\ln d_{E}=\ln\widetilde{d}_{1}(1+\frac{\ln\widetilde{d}_{2}}{\ln\widetilde{d}_{2}})=\ln\widetilde{d}_{1}(1+\frac{\ln D_{2}}{\ln D_{1}})=\ln\widetilde{d}_{1}\frac{\ln D}{\ln D_{1}}$. Rearranging the terms, we find, $\ln\widetilde{d}_{1}=\ln D_{1}(\frac{\ln d_{E}}{\ln D})$. Finally,
$\widetilde{d}_{1}=D_{1}^{\gamma},\mathrm{~{}with~{}}\gamma=\frac{\ln d_{E}}{\ln D}.$ (18)
This finding confirms our proposition that the result in Eq. 14 can be strengthened for a quantum system that obeys ETH. In other words, for a quantum system which thermalizes, the result of Page and Sen holds for the subspaces associated with each narrow energy window. We may here naively note that the limit $\gamma=1$ gives the results of Page and Sen.
Figure 3: The average VN entropy ($\overline{S}^{sb}_{VN}$) of subsystem is plotted against $\ln(DOS)$ across the energy spectrum (separately for the left and the right halves of the spectrum). Here the total system is described by the Hamiltonian of Eq. 20 and $\overline{S}^{sb}_{VN}$ is calculated as per Eq. 4.4. The calculations are done with the system (subsystem) size 16 (6) and $\Delta_{2}=0.5$.
To get numerical evidence to assert our proposition (Eq. 15), we first note that
$\overline{S}^{sb}_{VN}\approx\ln\widetilde{d}_{1}=(\frac{\ln D_{1}}{\ln D})\ln d_{E}=\lambda\ln d_{E},$ (19)
where $\lambda$ is the ratio of the size of the subsystem to the size of the total system (i.e., $\frac{l_{1}}{l_{1}+l_{2}}$). If we take unit energy interval, we find from Eq. 19 that $\overline{S}^{sb}_{VN}\propto\ln(DOS)$, where $DOS$ is the density-of-states of the full system at energy $E$. For numerical calculations, we take a spin-1/2 one-dimensional system with next-nearest neighbor interactions, as described by the Hamiltonian in Eq. 20. For each narrow energy interval, we evaluate $\overline{S}^{sb}_{VN}$ and $\ln(DOS)$ independently and plot them for different energies across the energy spectrum of the full system, as seen in Fig. 3. The linearity of the plot asserts our proposition that, for a system which obeys ETH, the result of Page and Sen holds for the subspaces associated with each narrow energy window.
Interestingly, the result presented above can also be interpreted as the statement of equivalence between the VN entropy and the TH entropy of a subsystem (of a larger system which thermalizes). We note that $\ln d_{E}$ is the TH entropy (more specifically, qBN entropy) of the full system at energy $E$. Assuming the additivity of the TH entropy, one can say that the TH entropy of the subsystem is proportional to $\ln d_{E}$. This finally shows, from Eq. 19, that the VN entropy and the TH entropy of a subsystem are equivalent.
@@
We now discuss the issue of equivalence using the concept of the canonical typicality. If the full system is in a pure state, for the overwhelming majority of pure states (of the total system), the subsystem can be described by a Gibbsian distribution $\rho_{\beta}$ [26, 11]. This important result immediately implies that, if the full system is in a pure state, for the overwhelming cases (i.e., for the typical states), the VN entropy of a subsystem will be equivalent to its TH entropy (qGB entropy), as in the case of an open system. The result of the canonical typicality leaves open the possibility of pure states (of the full system) for which a subsystem is not described by a Gibbsian distribution or for which a subsystem does not thermalize. We note that there are systems whose eigenstates (from a narrow energy window) are such that the eigen spectra of the corresponding reduced density matrices are quite different to each other. The VN entropy calculated from them has large deviation from the mean (see Fig. 2). These systems do not follow the eigenstate thermalization hypothesis or ETH [5, 33, 19]. For such systems, the VN entropy and the TH entropy (or qBN entropy) are not equivalent.
Figure 4: The subsystem VN entropy ($S^{sb}_{VN}$) when the full system is in a mixed state (see Eq. 4.4). Here, the system (subsystem) size is 16 (6).
4.4 Numerical results
In the part we take a spin system and numerically demonstrate the issue of equivalence between the VN entropy and the TH entropy of a subsystem in the different cases discussed earlier.